FEMTOSECOND CORRELATION SPECTROSCOPY

\underline{DONGPING ZHONG}, Departments of Physics, Chemistry and Biochemistry, 174 West 18th Avenue, The Ohio State University, Columbus, OH 43210; and AHMED H. ZEWAIL, Laboratory for Molecular Sciences, Arthur Amos Noyes Laboratory of Chemical Physics, California Institute of Technology, Pasadena, CA 91125.

We developed a kinetic-energy resolved time-of-flight mass spectrometry with femtosecond time resolution. The correlation of time, speed, orientation and state of reaction products was obtained. With these four resolution(s), a variety of chemical reactions was studied and new phenomena were discovered. Here, the classic example of bimolecular charge-transfer reactions of benzene with iodine and iodine monochloride will be given in detail.